OpenStudy (anonymous):

What is the distance each line: A) (-1,1), (0,4) B) (0,-4), (6,0) C) (6,0), (-1,1) The formula is √(x^2-x^1) (y^2-y^1) PLEASE HELP!!

OpenStudy (anonymous):

\[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (anonymous):

A = sqrt[ (4-1)^2 + (0--1)^2)] = sqrt 10 b = sqrt [(0--4)^2 + (6-0)^2 ] = sqrt 52 C = sqrt [ (1-0)^2 + (-1-6)^2] = sqrt 50

OpenStudy (anonymous):

a=\[\sqrt{10}\]

OpenStudy (anonymous):

THANK YOU TO ALL HELPED! YOU HAD BEEN REWARDED

OpenStudy (anonymous):

THANK YOU TO ALL HELPED! YOU HAD BEEN REWARDED